OpenStudy (horsegal244):

Steve set up a business making friendship bracelets. He earned $70 in January. He is projected to earn 120% more in February. How much is Steve projected to make in February?

OpenStudy (horsegal244):

@mathstudent55

OpenStudy (horsegal244):

@misty1212

OpenStudy (anonymous):

Multiply the $70 by 1.2 (equals 120%)

OpenStudy (anonymous):

looks good to me

OpenStudy (horsegal244):

ok.... I'm confused still

OpenStudy (anonymous):

actually it does not

OpenStudy (anonymous):

Its 120% more so after you add the 70 with the 70(1.2)

OpenStudy (anonymous):

to increase a number by \(120\%\) multiply it by \(100\%+120\%=1+1.2=2.2\)

OpenStudy (anonymous):

i.e. compute \[70\times 2.2\]

OpenStudy (horsegal244):

154

OpenStudy (horsegal244):

can u guys help me with a few more?

OpenStudy (anonymous):

Yes... I forgot there was the "more", not just 120%

OpenStudy (anonymous):

sure...

OpenStudy (horsegal244):

k